25 May 2010

Doing a lot of good deeds

tekstActivity in the gas laboratory. Thomas to the right and Remi in the middle are studying the behaviour of the instruments. The busy person is Vasileios.

The drillers report, that they soon are back on-line. Until then, several processers have participated in a number of little jobs that need to be done anyway.
In the CFA laboratory, people are satisfied. In the last 24 hours, the measurements have been running fine. Joe McConnell can now sit in a couch in the main dome and follow his black carbon measurements on-line on his computer. Also, the on-line water isotope measurements are running. We have now completed the moving of tents to the surface, and the last set of curtains is being mounted in the science trench.

What we have done today:

  1. Reaming and filtering the hole.
  2. Running in equipment in gas laboratory.
  3. Processing ice cores. Both main core and Joe McConnell’s Humbolt core were sampled.
  4. Moving Red Dome 1 to surface.
  5. Mounting curtains in front of core buffer.
  6. Measuring CFA. Today we measured 23.95 m of the SC piece. All systems online. The allocated SC-pieces have been measured, and the team now moves to the CFA pieces in the core storage.
  7. Water vapour sampling site now on-line.
  8. The on-line isotope measurements in the new warm laboratory are working.
  9. Making plastic sample bag inventory in camp.

Ad.1: Drillers report:
The reaming is now complete. Drillers then mounted the drill and drilled a short core both to clean the bottom of the hole and to make sure, there was no damage at the bottom due to the lost core barrels. They then proceeded to mount electronics section no.2 and the filter for a few filter runs. During these filter runs, electronics section no.2 failed completely, and the drillers now hope, that an intermittent error of section no.2 that has been plaguing section no.2 is permanent, so that it can tracked and be repaired. Drillers will mount section no.1 and continue drilling tomorrow.
Driller’s depth 1823.44 meters.  Logging depth 1836.95 meters

Ad.3: Processed 10 bags, last bag 3337, 1835.35 m. Cut 5 cm O-18 samples from 15 m Humbolt core.

Ad.5: The curtains already have an effect. Temperatures today in: drill trench -18°C, in tunnel -22°C, in core buffer – 20°C, in science trench -20°C and in ice cave -26°C. Cooling tunnels not yet switched on.

Weather: Overcast, in the evening clearing up . -13°C to - 7°C, 5-15 knots from SSE, later WSW. Visibility: Unrestricted or down to 1 mile during snow showers.

FL, J.P. Steffensen

  Floor is being laid for the red dome tent. The work was a little difficult due to snow. Christopher is hammering nails, Theo and Peter are carrying plywood, Sverrir knocks the sheets in place and Lou is nailing them.


← Previous entry   Next entry →